# The topology of a chaotic attractor in the Kuramoto-Sivashinsky equation

### v1

@@ III.3.2 Autoencoder return map
We proceed in the same way with the autoencoder’s latent coordinates $(h_{1},h_{2},h_{3})$ We define a Poincaré section $\mathcal{P}_{\mathrm{AE}}$ given by the plane $h_{2}=3/4$ and flow direction $\partial_{t}h_{2}>0$, so that the flow is approximately orthogonal to it. Figure 10 shows that this defines a unique crossing section along the band, giving us a sequence $\{h_{1}^{(i)}\}$. The return map $g_{1}$ is defined by $g_{1}:h_{1}^{(i)}\mapsto h_{1}^{(i+1)}$ (see top left panel of figure 11). We again fit the resulting smooth curve, with a polynomial of order three. We make the same observations as for the POD return map: $g_{1}$ seems to have one intersection with the identity $h_{1}^{(i)}=h_{1}^{(i+1)}$, implying one periodic orbit with $p=1$. Further iterations/compositions of the return map (see figure 11) indicate that there are 1, 0 and 1 periodic orbits with $p=2,3,4$ respectively. We again split the band domain $J=[0.198,0.247]$ at the minimum value of $g_{1}$, giving approximately $J_{0}=[0.198,0.23]$ and $J_{1}=[0.23,0.247]$. We observe similar qualitative dynamics to the POD return map: points starting in $J_{0}$ are mapped to $J$, while points starting in $J_{1}$ are mapped to a subset of $J_{0}$ (see figure fig. 12).
Figure 10: $(h_{1},h_{2})$ projection of the chaotic attractor in the latent space. The grey points are from a physical time-series that is put through the encoder. The magenta line illustrates the Poincaré section $h_{2}=3/4$ and $\partial_{t}h_{2}>0$, denoted $\mathcal{P}_{\mathrm{AE}}$.
@@
Figure 12: First return map for the Poincaré section $\mathcal{P}_{\mathrm{AE}}$. The blue dots show data from a chaotic timeseries, whereas the green dot, red line and black line show the periodic orbits $1_{1}$, $2_{1}$ and $4_{1}$ respectively (see table 1). Below the figure, the intervals represent the first and second level within the state space partition derived from the simplified return map shown in figure fig. 11. These intervals are used to construct symbolic names for the periodic orbits.
@@ III.4 Finding periodic orbits
We find periodic orbits (such as the one plotted in figures 3 and 6) by generating guesses using the methodology from Beck, Parker, and Schneider (2024) and converging them using the algorithms from Azimi, Ashtari, and Schneider (2022). More concretely, we obtain the guesses by defining arbitrary closed curves (loops) within the latent space, and decoding them to the physical space. This gives us a time-series that is already time-periodic and lies close on the chaotic attractor, however does not necessarily satisfy the PDE. The convergence algorithm then deforms the loop in an attempt to turn it into a solution, by minimizing a cost function $J$ that quantifies the misalignment between the tangent vectors and the flow vectors at each point of the loop. Thus, a root of the cost function corresponds to a closed curve in the state space that satisfies the flow equations everywhere, and is hence a periodic orbit.
@@
re 11 suggests that we have successfully identified all relevant periodic orbits with $p<7$. For example, we initially find periodic orbits with $p=3$, but upon visual inspection, we discarded all of those as not lying on the attractor. This agrees with our return maps. Since the only intersection of the third iterate with the identity (in the bottom left panel of figure 11) corresponds to the $p=1$ orbit, the iterated return map indeed suggests that no orbit with $p=3$ on the attractor exists. Likewise, all identified orbits exactly correspond to those expected from the return maps. For $p\geq 7$, it is difficult to be certain about the number of periodic orbits by purely looking at the return maps, and so we do not claim that we have found all of them with certainty. All orbits that appear to lie on the attractor were included in our analysis, and none were discarded purely on the basis of a return map argument. For the $15$ orbits, we arbitrarily chose names based on $p$ and a running index, as indicated in table 1.
$T$ $p$ Name
25.706 1 $1_{1}$
51.617 2 $2_{1}$
103.190 4 $4_{1}$
128.997 5 $5_{1}$
129.036 5 $5_{2}$
154.731 6 $6_{1}$
154.789 6 $6_{2}$
180.506 7 $7_{1}$
180.558 7 $7_{2}$
206.258 8 $8_{1}$
206.315 8 $8_{2}$
232.192 9 $9_{1}$
257.836 10 $10_{1}$
283.597 11 $11_{1}$
283.801 11 $11_{2}$
Table 1: Period $T$ and number of Poincaré intersections $p$ with the Poincaré sections $\mathcal{P}_{\mathrm{POD}}$ and $\mathcal{P}_{\mathrm{AE}}$ for each of the 15 periodic orbits found for the chaotic attractor of interest. The names are chosen based on $p$ and a running index in the order of increasing period $T$.
@@ IV Topological Analysis
The Birman-Wiliam theorem also states that periodic orbits of the strange attractor are in one-to-one correspondence with the periodic orbits on the template, with at most two extraneous orbits(Birman and Williams, 1983b; Ghrist, 1997). Moreover, the template has the property that all periodic orbits can be projected onto it without altering any of their topological invariants, crucially including their (self-)linking numbers. Certain preconditions of the Birman-Wiliam theorem can be relaxed; see Chapter 13 in Gilmore and Lefranc (2003) for more details. Given an higher dimensional flow whose Lyapunov dimension is less than 3, one can first project the flow along its stable directions to an intermediate manifold where we can apply the Birman-Wiliam theorem. In the following, we attempt to construct the template of the chaotic attractor from the identified set of periodic orbits and thereby identify the structure of the spectrum of periodic orbits on the attractor of the PDE.
@@ IV.1 Method
A template provides a way of enumerating the periodic orbits by assigning a unique symbol to each branch, yielding a descriptive name based on the symbolic sequence that the orbit traverses the branches(Plumecoq and Lefranc, 1999a; Gilmore and Lefranc, 2003). Let $A=(a_{1},\dots,a_{p_{A}})$ and $B=(b_{1},\dots,b_{p_{B}})$ be two periodic orbits written with their symbolic names. Here, $p_{A}$ and $p_{B}$ denote the number of intersection points of the respective orbits with a Poincaré section. The relative rotation rates $R_{i,j}(A,B)$ describe how much, on average, two orbits $A$ and $B$ rotate around one another when starting from initial conditions $a_{i}$ and $b_{j}$(Tufillaro, Solari, and Gilmore, 1990).
@@
In this way, the linking numbers can be expressed as polynomials in the torsion and layering terms, with coefficients depending on the symbolic name assigned to each orbit (see Appendix A.2.5 in Gilmore and Lefranc (2003)),
@@
  • Iterate over the possible names for the orbits (and/or the the number of branches) and solve the system of equations (7). If there is a unique valid solution one can include additional orbits to check the validity of a candidate. If, on the other hand, multiple valid names and template candidates are found, the problem is undetermined and more orbits must be included.
Hence the algorithm returns valid set(s) of possible names together with a template. The procedure is described precisely in Appendix A in Gilmore and Lefranc (2003). An overview of the topological analysis program is given in Lefranc (2007). The previous algorithm does not immediately extend to flows in dimensions higher than three. However, in the considered case of the KSE, the dynamic can be projected in $\mathbb{R}^{3}$ where the computations of linking numbers can be carried out.
@@ IV.2 A template for the KSE
Strange attractors can be classified by the topological organization of periodic orbitsLetellier, Stankevich, and Rössler (2022). If the dynamic can be embedded in three dimension then the periodic orbits form knots and links and their topological organization is described by their linking numbers. Then, by the Birman-Wiliam theorem, one can project the flow onto a template while preserving the organization of the periodic orbits. A template not only supports all the periodic orbits and describes their topological organization but also offers a natural way for enumerating these periodic orbits. For $L\approx 18.05$, the Lyapunov dimension associated to the KSE system is lower than three. We use two distinct tools: proper orthogonal decomposition and an autoencoder neural network, described earlier, to embed the dynamic in 3D latent spaces. Then we compare the templates resulting from each method.
@@ IV.2.1 POD template
Figure 13: Template of KSE system with $L\approx 18.05$ extracted from linking numbers calculated on POD projections of UPOs. This represents the dynamics of the system in a topological manner, as trajectories move clockwise around this figure. All UPOs must follow the branched manifold with labels of the two branches yielding symbolic names for the orbits. Two low-period orbits used to identify this template are shown in red ($1_{1}$) and blue ($2_{1}$).
@@
After projecting the periodic orbits found in Section III.4 onto the POD space we compute their linking numbers using the algorithm described by Qu and James (2021). The linking numbers are given in table 2. We did not calculate self-linking numbers for the periodic orbits, as this requires the choice of a framing, and it is not clear how to do this consistently given the dimensionality reduction. It is not necessary to compute self-linking numbers to determine a template, but without them it requires more periodic orbits to find a unique template(Rosalie, 2016).
Some of the UPOs considered have very long periods. We found that numerically calculating the linking numbers of these requires discretisation of each loop using a very high number of points, given the necessity of accurately counting the crossings of periodic orbits in very close proximity. To resolve the most challenging cases, it was necessary to discretise the orbits with $2^{13}$ points.
The first-return map shown in figure 8 has two monotonic branches and so the template has two branches labelled 0 and 1. With the five lowest period UPOs as input to the algorithm described in section IV.1, we obtain a unique candidate template with two branches. The candidate template is represented in figure 13 and is uniquely characterised by the following algebraic description:
$(t_{0,0},t_{0,1},t_{1,1},l_{0,1})=(-1,-2,-2,1).$ (8)
@@ IV.2.2 Autoencoder template
Figure 14: As for fig. 13 but with linking numbers calculated on autoencoder latent space projections of UPOs.
@@
Using instead a projection into the latent space of the autoencoder, the linking numbers were identical except for a global change of sign, i.e. all negative linking numbers in table 2 were replaced by positive numbers. Again, the first-return map shown in figure 12 has two monotonic branches and so the template has two branches labelled 0 and 1. With the five lowest period UPOs as input to the algorithm, we obtain a unique candidate template with two branches represented in figure 14 whose algebraic description is given by:
$(t_{0,0},t_{0,1},t_{1,1},l_{0,1})=(1,2,2,-1).$ (9)
@@
The next ten orbits are incorporated into the input set, leading to the addition of 90 equations to the system (7). All those are consistent with the template found. The resulting template is the mirror-image of the one obtained with the POD method. By taking the template associated with the POD represented in figure 13, we can flip it (so that the ”bottom” becomes the ”front”). Then, by taking the mirror image, we obtain the template associated with the autoencoder in figure 14. The two templates are diffeomorphic, but not isotopic, see Section 12 in Gilmore and Lefranc (2003). The topological organisation of the periodic orbits is preserved up to a global change of sign. Consequently, the characterisation is independent, as intended, of the different coordinates resulting from different dimensionality-reduction techniques.
IV.3 Symbolic names
@@
A template provides a natural way to enumerate the periodic orbits: to each periodic orbit, one assigns a word whose letters represent the successive branches traversed. A collection of periodic orbits with unique symbolic names are a key element of a method for constructing a generating partition by interpolating the encoding map (Plumecoq and Lefranc, 1999a, b).
@@
During the template identification we obtained several sets of possible names so, according to the algorithm described in section IV.A, we should add more orbits in the input set. However, in our particular case of the KSE, we are able to determine the symbolic names of the input set of periodic orbits using the POD return map as shown in fig. 9. The set of symbolic names for the orbits in table 1 is presented in table 3.
$1_{1}$ $(0)$ $5_{2}$ $(00101)$ $7_{2}$ $(1000010)$ $10_{1}$ $(0000000101)$
$2_{1}$ $(10)$ $6_{1}$ $(100000)$ $8_{1}$ $(01000000)$ $11_{1}$ $(01010000000)$
$4_{1}$ $(0001)$ $6_{2}$ $(000101)$ $8_{2}$ $(10100000)$ $11_{2}$ $(00001010101)$
$5_{1}$ $(10000)$ $7_{1}$ $(1000000)$ $9_{1}$ $(100001010)$
Table 3: Set of possible names for the first 15 orbits of the KSE system with $L\approx 18.05$
@@
To check the validity of this set of symbolic names, we compute the expression of the linking numbers in terms of layering and torsion coefficients with respect to the symbolic names of the periodic orbits (right side in (7)). Then, we substitute the coefficients with the template identification found previously (8). In all cases, there is agreement with the linking numbers obtained from the embedded periodic orbits in the POD space in table 2 (left handside in (7)). For instance, consider
$Lk(000101,00001010101)$ $=l_{01}(\pi_{0}-\pi_{1}+12)$
$+14t_{0,0}+15t_{01}+4t_{11}$
$Lk(0001,00101)$ $=\frac{l_{01}}{2}(-4\pi_{0}\pi_{1}+\pi_{0}+2\pi_{1}+6)$
$+\frac{9}{2}t_{00}+\frac{9}{2}t_{01}+t_{11}.$
@@
$Lk(000101,00001010101)$ $=-39=Lk(6_{2},11_{2})$
$Lk(0001,00101)$ $=-12=Lk(4_{1},5_{2}).$
We proceed similarly with the autoencoder method. Again, as shown in figure 12, from the autoencoder return map we obtain the same set of symbolic words for the input set of periodic orbits. The linking numbers obtained from the symbolic names and the template identification (9) are consistent with the linking numbers obtained from the embedded periodic orbits in the autoencoder latent space.
Thus consistent symbolic names for low-period UPOs are derived from two different dimensionality reduction methods and the symbolic names are validated by the template identification.

### v2

@@ III.3.2 Autoencoder return map
We proceed in the same way with the autoencoder’s latent coordinates $(h_{1},h_{2},h_{3})$ We define a Poincaré section $\mathcal{P}_{\mathrm{AE}}$ given by the plane $h_{2}=3/4$ and flow direction $\partial_{t}h_{2}>0$, so that the flow is approximately orthogonal to it. Figure 10 shows that this defines a unique crossing section along the band, giving us a sequence $\{h_{1}^{(i)}\}$. The return map $g_{1}$ is defined by $g_{1}:h_{1}^{(i)}\mapsto h_{1}^{(i+1)}$ (see top left panel of figure 11). We again fit the resulting smooth curve, with a polynomial of order three. We make the same observations as for the POD return map: $g_{1}$ seems to have one intersection with the identity $h_{1}^{(i)}=h_{1}^{(i+1)}$, implying one periodic orbit with $p=1$. Further iterations/compositions of the return map (see figure 11) indicate that there are 1, 0 and 1 periodic orbits with $p=2,3,4$ respectively. We again split the band domain $J=[0.198,0.247]$ at the minimum value of $g_{1}$, giving approximately $J_{1}=[0.198,0.23]$ and $J_{2}=[0.23,0.247]$. We observe similar qualitative dynamics to the POD return map: points starting in $J_{1}$ are mapped to $J$, while points starting in $J_{2}$ are mapped to a subset of $J_{1}$ (see figure 12).
Figure 10: $(h_{1},h_{2})$ image of the chaotic attractor in the latent space of the autoencoder. The grey points are from a physical timeseries that is put through the encoder. The magenta line illustrates the Poincaré section $h_{2}=3/4$ and $\partial_{t}h_{2}>0$, denoted $\mathcal{P}_{\mathrm{AE}}$.
@@
Figure 12: First return map for the Poincaré section $\mathcal{P}_{\mathrm{AE}}$. The blue dots show data from a chaotic timeseries, whereas the green dot, red line and black line show the periodic orbits $1_{1}$, $2_{1}$ and $4_{1}$ respectively (see table 1). Below the figure, the intervals represent the first and second level within the state space partition derived from the simplified return map shown in figure 11. These intervals are used to construct symbolic sequences for the periodic orbits.
@@ III.4 Finding periodic orbits
We find periodic orbits (such as the one plotted in figures 3 and 6) by generating guesses using the methodology from Beck, Parker, and Schneider (2024) and converging them using the algorithms from Azimi, Ashtari, and Schneider (2022). This is similar to the variational method used by Dong (2018) to find UPOs and symbolic dynamics at a higher $L$ value in the KSE. We obtain the guesses by defining arbitrary closed curves (loops) within the latent space, and decoding them to the physical space. This gives us a timeseries that is already time-periodic and lies close on the chaotic attractor, however does not necessarily satisfy the PDE. The convergence algorithm then deforms the loop in an attempt to turn it into a solution, by minimizing a cost function $J$ that quantifies the misalignment between the tangent vectors and the flow vectors at each point of the loop. Thus, a root of the cost function corresponds to a closed curve in the state space that satisfies the flow equations everywhere, and is hence a periodic orbit.
@@
Comparing the extracted list of 15 periodic orbits with the return map and its iterates in figure 11 suggests that we have successfully identified all relevant periodic orbits with $p<7$. For example, we initially find periodic orbits with $p=3$, but upon visual inspection, we discarded all of those as not lying on the attractor. This agrees with our return maps. Since the only intersection of the third iterate with the identity (in the bottom left panel of figure 11) corresponds to the $p=1$ orbit, the iterated return map indeed suggests that no orbit with $p=3$ on the attractor exists. This is consistent with the symbolic dynamics derived from the approximate return map as shown in figure 12, since period 3 symbolic sequences are not admissible (Hao and Zheng, 1998). Recall that the existence of a period 3 orbit implies chaos, but the converse is not true. Likewise, all identified orbits exactly correspond to those expected from the return maps. For $p\geq 7$, it is difficult to be certain about the number of periodic orbits by purely looking at the return maps, and so we do not claim that we have found all of them with certainty. All orbits that appear to lie on the attractor were included in our analysis, and none were discarded purely on the basis of a return map argument. For the $15$ orbits, we arbitrarily chose names based on $p$ and a running index, as indicated in table 1.
$T$ $p$ Name Symbolic sequence
25.706 1 $1_{1}$ $(1)$
51.617 2 $2_{1}$ $(21)$
103.190 4 $4_{1}$ $(2111)$
128.997 5 $5_{1}$ $(21111)$
129.036 5 $5_{2}$ $(21121)$
154.731 6 $6_{1}$ $(211111)$
154.789 6 $6_{2}$ $(211121)$
180.506 7 $7_{1}$ $(2111111)$
180.558 7 $7_{2}$ $(2111121)$
206.258 8 $8_{1}$ $(21111111)$
206.315 8 $8_{2}$ $(21111121)$
232.192 9 $9_{1}$ $(211112121)$
257.836 10 $10_{1}$ $(2111111121)$
283.597 11 $11_{1}$ $(21111111121)$
283.801 11 $11_{2}$ $(21111212121)$
Table 1: Period $T$ and number of Poincaré intersections $p$ with the Poincaré sections $\mathcal{P}_{\mathrm{POD}}$ and $\mathcal{P}_{\mathrm{AE}}$ for each of the 15 periodic orbits found for the chaotic attractor of interest. The names are chosen based on $p$ and a running index in the order of increasing period $T$. Two identical set of symbolic sequences is obtained from the first return maps for their respective Poincaré sections $\mathcal{P}_{\mathrm{POD}}$ and $\mathcal{P}_{\mathrm{AE}}$.
@@ IV Topological Analysis
The Birman-Wiliam theorem also states that periodic orbits of the strange attractor are in one-to-one correspondence with the periodic orbits on the template, with at most two extraneous orbits(Birman and Williams, 1983b; Ghrist, 1997). Moreover, the template has the property that all periodic orbits can be projected onto it without altering any of their topological invariants, crucially including their (self-)linking numbers. Certain preconditions of the Birman-Wiliam theorem can be relaxed; see Chapter 13 in Gilmore and Lefranc (2003) for more details. Given a higher dimensional flow whose Lyapunov dimension is less than 3, one can first project the flow along its stable directions to an intermediate manifold where we can apply the Birman-Wiliam theorem. In the following, we attempt to construct the template of the chaotic attractor from the identified set of periodic orbits and thereby identify the structure of the spectrum of periodic orbits on the attractor of the PDE.
@@ IV.1 Method
A template provides a way of enumerating the periodic orbits by assigning a unique symbol to each branch, yielding a descriptive name based on the symbolic sequence that the orbit traverses the branches(Plumecoq and Lefranc, 1999a; Gilmore and Lefranc, 2003). Let $A=(a_{1},\dots,a_{p_{A}})$ and $B=(b_{1},\dots,b_{p_{B}})$ be two periodic orbits written with their symbolic sequences. Here, $p_{A}$ and $p_{B}$ denote the number of intersection points of the respective orbits with a Poincaré section. The relative rotation rates $R_{i,j}(A,B)$ describe how much, on average, two orbits $A$ and $B$ rotate around one another when starting from initial conditions $a_{i}$ and $b_{j}$(Tufillaro, Solari, and Gilmore, 1990).
@@
In this way, the linking numbers can be expressed as polynomials in the torsion and layering terms, with coefficients depending on the symbolic sequence assigned to each orbit (see Appendix A.2.5 in Gilmore and Lefranc (2003)),
@@
  • Iterate over the possible symbolic sequences for the orbits (and/or the the number of branches) and solve the system of equations (7). If there is a unique valid solution one can include additional orbits to check the validity of a candidate. If, on the other hand, multiple valid names and template candidates are found, the problem is undetermined and more orbits must be included.
Hence the algorithm returns valid set(s) of possible names together with a template. The procedure is described precisely in Appendix A in Gilmore and Lefranc (2003). An overview of the topological analysis program is given in Lefranc (2007). The previous algorithm does not immediately extend to flows in dimensions higher than three. However, in the considered case of the KSE, it appears that the attractor can be embedded in $\mathbb{R}^{3}$ where the computations of linking numbers can be carried out.
@@ IV.2 A template for the KSE
Strange attractors can be classified by the topological organization of periodic orbitsLetellier, Stankevich, and Rössler (2022). If the dynamics can be embedded in three dimension then the periodic orbits form knots and links and their topological organization is described by their linking numbers. Then, by the Birman-Wiliam theorem, one can project the flow onto a template while preserving the organization of the periodic orbits. A template not only supports all the periodic orbits and describes their topological organization but also offers a natural way for enumerating these periodic orbits. For $L\approx 18.05$, the Lyapunov dimension associated to the KSE system is lower than three. We use two distinct tools: proper orthogonal decomposition and an autoencoder neural network, described earlier, to embed the dynamic in 3D latent spaces. Then we compare the templates resulting from each method.
@@ IV.2.1 POD template
Figure 13: Template of KSE system with $L\approx 18.05$ extracted from linking numbers calculated on POD projections of UPOs. This represents the dynamics of the system in a topological manner, as trajectories move clockwise around this figure. All UPOs must follow the branched manifold with labels of the two branches yielding symbolic sequences for the orbits. Two low-period orbits used to identify this template are shown in red $(1)$ and blue $(21)$
@@
The simplified first-return map shown in figure 8 has two monotonic branches and so the template has two branches labelled 1 and 2. With the five lowest period UPOs as input to the algorithm described in section IV.1, we obtain a unique candidate template with two branches. The candidate template is represented in figure 13 and is uniquely characterised by the following algebraic description:
$(t_{1,1},t_{1,2},t_{2,2},l_{1,2})=(-1,-2,-2,1).$ (8)
@@
Note that after projecting the periodic orbits found in Section III.4 onto the POD space we compute their linking numbers using the algorithm described by Qu and James (2021). The linking numbers are given in table 2. We did not calculate self-linking numbers for the periodic orbits, as this requires the choice of a framing, and it is not clear how to do this consistently given the dimensionality reduction. It is not necessary to compute self-linking numbers to determine a template, but without them it requires more periodic orbits to find a unique template(Rosalie, 2016).
Some of the UPOs considered have very long periods. We found that numerically calculating the linking numbers of these requires discretization of each loop using a very high number of points, given the necessity of accurately counting the crossings of periodic orbits in very close proximity. To resolve the most challenging cases, it was necessary to discretize the orbits with $2^{13}$ points.
@@ IV.2.2 Autoencoder template
Figure 14: As for figure 13 but with linking numbers calculated on autoencoder latent space images of UPOs.
@@
Using instead the images of the UPOs in the latent space of the autoencoder, the linking numbers were identical except for a global change of sign, i.e. all negative linking numbers in table 2 were replaced by positive numbers. Again, the first-return map shown in figure 12 has two monotonic branches and so the template has two branches labelled 1 and 2. With the five lowest period UPOs as input to the algorithm, we obtain a unique candidate template with two branches represented in figure 14 whose algebraic description is given by:
$(t_{1,1},t_{1,2},t_{2,2},l_{1,2})=(1,2,2,-1).$ (9)
@@
The next ten orbits are incorporated into the input set, leading to the addition of 90 equations to the system (7). All those are consistent with the template found. The resulting template is the mirror-image of the one obtained with the POD method. By taking the template associated with the POD represented in figure 13, we can flip it (so that the “bottom” becomes the “front”). Then, by taking the mirror image, we obtain the template associated with the autoencoder in figure 14. The two templates are diffeomorphic, but not isotopic(Gilmore, 2013). The topological organisation of the periodic orbits is preserved up to a global change of sign. Consequently, the characterisation is independent, as intended, of the different coordinates resulting from different dimensionality-reduction techniques.
IV.3 Symbolic sequences
@@
A template provides a natural way to enumerate the periodic orbits: to each periodic orbit, one assigns a word whose letters represent the successive branches traversed. A collection of periodic orbits with unique symbolic sequences are a key element of a method for constructing a generating partition by interpolating the encoding map (Plumecoq and Lefranc, 1999a, b).
@@
During the template identification we obtained several sets of possible names so, according to the algorithm described in section IV.A, we should add more orbits in the input set. However, in our particular case of the KSE, we are able to determine the symbolic sequences of the input set of periodic orbits using the POD return map as shown in figure 9. The set of symbolic sequences for the input set of orbits is presented in table 1. The choice of the starting location of a symbolic sequence for a given orbit is arbitrary; we follow the convention given by Hao and Zheng (1998), i.e. choosing the sequence whose initial string has even parity.
@@
To check the validity of this set of symbolic sequences, we compute the expression of the linking numbers in terms of layering and torsion coefficients with respect to the symbolic sequences of the periodic orbits (right side in (7)). Then, we substitute the coefficients with the template identification found previously (8). In all cases, there is agreement with the linking numbers obtained from the embedded periodic orbits in the POD space in table 2 (left handside in (7)). For instance, consider
$Lk(211121,21111212121)$ $=l_{12}(\pi_{1}-\pi_{2}+12)$
$+14t_{11}+15t_{12}+4t_{22}$
$Lk(2111,21121)$ $=\frac{l_{12}}{2}(-4\pi_{1}\pi_{2}+\pi_{1}+2\pi_{2}+6)$
$+\frac{9}{2}t_{11}+\frac{9}{2}t_{12}+t_{22}.$
@@
$Lk(211121,21111212121)$ $=-39=Lk(6_{2},11_{2})$
$Lk(2111,21121)$ $=-12=Lk(4_{1},5_{2}).$
We proceed similarly with the autoencoder method. Again, as shown in figure 12, from the autoencoder return map we obtain the same set of symbolic words for the input set of periodic orbits. The linking numbers obtained from the symbolic sequences and the template identification (9) are consistent with the linking numbers obtained from the embedded periodic orbits in the autoencoder latent space.
Thus consistent symbolic sequences for low-period UPOs are derived from two different dimensionality reduction methods and the symbolic sequences are validated by the template identification.